# FlexControl: Computation-Aware ControlNet with Differentiable Router for Text-to-Image Generation

### v1

@@ Abstract
ControlNet offers a powerful way to guide diffusion‐based generative models, yet most implementations rely on ad-hoc heuristics to choose which network blocks to control—an approach that varies unpredictably with different tasks. To address this gap, we propose FlexControl, a novel framework that copies all diffusion blocks during training and employs a trainable gating mechanism to dynamically select which blocks to activate at each denoising step. With introducing a computation-aware loss, we can encourage control blocks only to activate when it benefit the generation quality. By eliminating manual block selection, FlexControl enhances adaptability across diverse tasks and streamlines the design pipeline, with computation-aware training loss in an end-to-end training manner. Through comprehensive experiments on both UNet (e.g., SD1.5) and DiT (e.g., SD3.0), we show that our method outperforms existing ControlNet variants in certain key aspects of interest. As evidenced by both quantitative and qualitative evaluations, FlexControl preserves or enhances image fidelity while also reducing computational overhead by selectively activating the most relevant blocks. These results underscore the potential of a flexible, data‐driven approach for controlled diffusion and open new avenues for efficient generative model design.
Figure 1: Dynamically control image generation based on timestep and specific sample. (a) The architecture of ControlNet. (b) The architecture of proposed FlexControl. (c) Statistics on the number of activated control blocks of FlexControl at each denoising step. Here, “50% spasity” indicates that the number of floating-point operations (FLOPs) of activated blocks is limited to 50% of the trainable branch.
@@ 3. Universal plug-and-play integration:
Our method seamlessly integrates with any dual-stream control-architecture, introducing minimal additional parameters and zero architectural modifications to host models. It enables flexible switching between full control and efficiency-optimized modes, depending on computational requirements.
@@ 2 Related Work
Efforts to enhance ControlNet’s efficacy and efficiency have focused on architectural redesigns, training optimizations, and inference acceleration. ControlNeXt [35] replaces ControlNet’s bulky auxiliary branches with a streamlined architecture and substitutes zero convolutions with Cross Normalization, slashing learnable parameters by 90% while maintaining stable convergence. Beyond architecture, multi-expert diffusion frameworks [25, 55] tailor denoising operations to specific timesteps, though their computational demands hinder practicality. To reduce inference costs, pruning techniques[12, 23, 13] trim redundant parameters from pre-trained denoising models, while distillation methods[16] train lightweight guide models to minimize denoising steps. Inspired by RepVGG [8], RepControlNet [6] introduces a novel reparameterization strategy: modal-specific adapters modulate features during training, and their weights are later merged with the base network, eliminating auxiliary computations at inference. Unlike prior methods that rely on fixed heuristics, post-hoc pruning, or static architectural modifications, our FlexControl introduces a dynamic, end-to-end trainable framework where block activation is both task-aware and computation-aware. By integrating a gating mechanism with a computational efficiency objective, our approach uniquely balances precision and resource usage, enabling adaptive control across diverse architectures (UNet, DiT) without manual intervention — a paradigm shift from rigid, task-specific designs to flexible, generalizable control.
@@ 3 Methodology
Figure 2: Overview of dynamic routing guided by router unit. (a) In the training stage, Gumbel noise is added to the discrete mask to assist the gradient backpropagation. (b) In the inference stage, the router unit controls whether to activate the control block according to the input latent variable. Once output the instruction of inactive, the corresponding control block and zero module will be skipped.
@@ 4.1 Quantitative comparison
Table 3: Complexity comparison on SD1.5 and SD3.0. We compare model parameters, FLOPs, and inference speed (it/s, iterations per second). The best values are highlighted in red, while the second-best values are shown in blue. FlexControl significant reduce overall FLOPs and inference time from ControlNet-Large.
@@ 4.3 Ablation study
Table 4: Quantitative comparison with existing methods on SD1.5. We report FID (↓), CLIP score (↑), and mIoU (↑) across different models. The best values and second-best are highlighted in red and blue. FlexControl outperform original ControlNet with only 0.3 activation blocks, while increasing the blocks budgets observed performance increasing. Noticeable, ControlNet-Large activate all blocks yet not out-perform our methods, highlight effective of our dynamic strategy.
@@
Table 5: Quantitative comparison with existing methods on SD3.0. We report FID (↓), CLIP score (↑), and SSIM (↑). The best and second-best values are highlighted in red and blue.FlexControl outperform original ControlNet with only 0.3 activation blocks, while increasing the blocks budgets observed performance increasing even more significant improvement than we observed in SD1.5.

### v2

@@ Abstract
ControlNet offers a powerful way to guide diffusion‐based generative models, yet most implementations rely on ad-hoc heuristics to choose which network blocks to control—an approach that varies unpredictably with different tasks. To address this gap, we propose FlexControl, a novel framework that copies all diffusion blocks during training and employs a trainable gating mechanism to dynamically select which blocks to activate at each denoising step. With introducing a computation-aware loss, we can encourage control blocks only to activate when it benefit the generation quality. By eliminating manual block selection, FlexControl enhances adaptability across diverse tasks and streamlines the design pipeline, with computation-aware training loss in an end-to-end training manner. Through comprehensive experiments on both UNet (e.g., SD1.5) and DiT (e.g., SD3.0), we show that our method outperforms existing ControlNet variants in certain key aspects of interest. As evidenced by both quantitative and qualitative evaluations, FlexControl preserves or enhances image fidelity while also reducing computational overhead by selectively activating the most relevant blocks. These results underscore the potential of a flexible, data‐driven approach for controlled diffusion and open new avenues for efficient generative model design. The code will soon be available at https://github.com/Anonymousuuser/FlexControl.
Figure 1: Dynamically inject conditional controls for image generation based on timestep and specific sample. (a) The architecture of the ControlNet. (b) The architecture of the proposed FlexControl. (c) Statistics on the number of activated control blocks of the FlexControl at each denoising step. Here, “50% spasity” indicates that the number of floating-point operations (FLOPs) of activated blocks is limited to 50% of the trainable branch.
@@ 3. Universal plug-and-play integration:
Our method seamlessly integrates with any dual-stream control-model, introducing minimal additional parameters and zero architectural modifications to host models. It enables flexible switching between full control and efficiency-optimized modes, depending on computational requirements.
@@ 2 Related Work
Efforts to enhance ControlNet’s efficacy and efficiency have focused on architectural redesigns, training optimizations, and inference acceleration. ControlNeXt [35] replaces ControlNet’s bulky auxiliary branches with a streamlined architecture and substitutes zero modules with Cross Normalization, slashing learnable parameters by 90% while maintaining stable convergence. Beyond this, multi-expert diffusion frameworks [25, 55] tailor denoising operations to specific timesteps, though their computational demands hinder practicality. To reduce inference costs, pruning techniques[12, 23, 13] trim redundant parameters from pre-trained denoising models, while distillation methods[16] train lightweight guide models to minimize denoising steps. Inspired by RepVGG [8], RepControlNet [6] introduces a novel reparameterization strategy: modal-specific adapters modulate features during training, and their weights are later merged with the base network, eliminating auxiliary computations at inference. Unlike prior methods that rely on fixed heuristics, post-hoc pruning, or static architectural modifications, our FlexControl introduces a dynamic, end-to-end trainable framework where block activation is both task-aware and computation-aware. By integrating a gating mechanism with a computational efficiency objective, our approach uniquely balances precision and resource usage, enabling adaptive control across diverse architectures (UNet, DiT) without manual intervention — a paradigm shift from rigid, task-specific designs to flexible, generalizable control.
@@ 3 Methodology
Figure 2: Overview of dynamic routing guided by the router unit. (a) In the training stage, Gumbel noise is added to the discrete mask to assist the gradient backpropagation. (b) In the inference stage, the router unit controls whether to activate the control block and whether to inject conditional control into the frozen block of the backbone according to the input latent variable. Once output the instruction of inactive, the corresponding control block and zero module will be skipped.
@@ 4.1 Quantitative comparison
Table 3: Complexity comparison on SD1.5 and SD3.0. We compare model parameters, FLOPs, and inference speed (it/s (↑), iterations per second and s/it (↓), seconds per iteration). The best values are highlighted in red, while the second-best values are shown in blue. FlexControl significant reduce overall FLOPs and inference time from ControlNet-Large.
@@ 4.2 Qualitative comparison
Table 4: Quantitative comparison with existing methods on SD1.5. We report FID (↓), CLIP_score (↑), and mIoU (↑) across different models. The best values and second-best are highlighted in red and blue. FlexControl outperform original ControlNet with less computation, while increasing the blocks budgets observed performance increasing. Noticeable, ControlNet-Large activate all blocks yet not out-perform our methods, highlight effective of our dynamic strategy.
4.3 Ablation study
@@
Table 5: Quantitative comparison with existing methods on SD3.0. We report FID (↓), CLIP_score (↑), and SSIM (↑). The best and second-best values are highlighted in red and blue. FlexControl outperform original ControlNet with less computation, while increasing the blocks budgets observed performance increasing even more significant improvement than we observed in SD1.5.
Results on DiT-based model.